OpenStudy (mitu12):

Given where (h,k)=(2,-3), answer the following questions. y=f(x-h)^2+k 1. What is the parent function?-Quadratic 2. Is the 3 translation vertical or horizontal? 3. Is the 2 translation vertical or horizontal?

OpenStudy (mitu12):

@mathstudent55

OpenStudy (mitu12):

@pooja195

OpenStudy (mitu12):

@sleepyjess

OpenStudy (mathmale):

Yes, the parent function is the quadratic function y=x^2. How would the graph of y=x^2 change if you were to graph y=(x-2)^2? Try graphing this. How would the graph of y=x^2 change if you were to graph y=(x)^2+3? Graph this. Now attempt to answer the original question. For more info, please review "vertical and horizontal translations" in your textbook or online learning materials.